OpenStudy (juanredondo2000):

Solve -2(2x + 5) - 3 = -3(x - 1).

OpenStudy (agreene):

start by distributing the constants: \[\color{red}{-2}(2x+5)-3=\color{red}{-3}(x-1)\]

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\) After you do what agreene asked, You'll have: \(\color{red}{-4x}-10-3=\color{red}{3x}+3\) Combine like terms is the next step. \(-4x\color{blue}{-13}=3x+3\) Now you have variables on both sides. \(-4x-13=3x+3\) Can you solve it from there?